OpenStudy (anonymous):

The value of Mr. Cox’s car x years after its purchase is given by the function, V(x) = 20,000(0.85)x. Approximately, what was the value of Mr. Cox’s car 5 years after its purchase?

OpenStudy (anonymous):

can you help me please!!!!!!!

OpenStudy (anonymous):

I dont understand how to get the answer for ^^^^^^^^ problem...... Similar Problem: The value of Mr. Cox's car x years after its purchase is given by the function, V(x) = 30,000(0.90)x. Approximately, what was the value of Mr. Cox's car 5 years after its purchase? V(x) = 30,000(0.90)x V(x) = 30,000(0.90)5 V(x) = 30,000(0.59049) V(x) = 177147

OpenStudy (anonymous):

It should look like this: 20,000(0.85)(5) Just multiply across and you'll get the answer! X represents the number of years so you switch 5 with x and you're good to go

OpenStudy (anonymous):

20,000x0.85x5?

OpenStudy (anonymous):

@chrisisbad92

OpenStudy (anonymous):

Yup, go for it.

OpenStudy (anonymous):

ok can you help me with another one ?

OpenStudy (anonymous):

Sure thing.